# Examining Racial Stereotypes in YouTube Autocomplete Suggestions

### v2

@@ 4.5 Manner
Inappropriate Behavior. This category contains autocompletes that describe Black people as lazy or behaving inappropriately in social settings. One longstanding negative stereotype of Black people maintains that they are lazy, which is also posed as an explanation for other undesirable attributes, such as their disproportionate poverty (Gilens, 1996). The autocomplete “black people can’t hear smoke detectors” depicts this stereotype; it led to videos that show Black people as being too lazy to change the low battery of smoke detectors even though their alarms kept sounding. Figure 9 presents a series of YouTube shorts that show how creators promote this stereotype.
@@ 5 Discussion
We began this study with the goal of examining the information that YouTube’s search feature provides to users seeking video content on race-related topics. We found that besides listing offerings that assuage users’ curiosities on these topics, YouTube plays a role in spreading racial stereotypes through its autocomplete function, a feature that incidentally cannot be turned off.
Our analysis reveals that race-related autocompletes embody troubling stereotypes regarding Blacks’ personal hygiene, financial abilities, language use, aggression, and laziness; Whites’ athletic abilities; and Asians’ image as the high-achieving model minority. Notably, we observed that YouTube search amplifies distorted racial views that appear in mainstream entertainment media by reproducing titles or dialogues from existing media content (e.g., “white men can’t jump”). We posit that this can be attributed to YouTube’s participatory culture (Khan and Malik, 2022), where users actively contribute to content creation aimed at capturing popular attention and resulting in monetary gains. Thus, an interplay of mainstream media and amateur online cultures could be perpetuating racial stereotypes on the platform. Creators’ communal and commercial aspirations (Boxman-Shabtai, 2019) could also account for our finding that racial stereotypes on YouTube are often subtly delivered in the form of ethnic humor and memes. These genres and formats may make naive viewers embrace group stereotyping as a new and enjoyable form of entertainment rather than recognize its potential harm toward racial minorities, as evidenced by the millions of views and likes such content generates.
In addition to reinforcing racial stereotypes, our analysis of YouTube autocompletes reflects tensions about interracial relationships, and Blacks’ and Hispanics’ concerns about their media representation. The norms of whiteness appear to prevail, with racial minorities struggling to combat structural inequalities (“can a black man be captain america,” “is there a hispanic disney princess”). These results serve as a lens through which to examine the current state of race in the U.S. Though some popular narratives maintain that we live in a ‘post-race’ society (Squires, 2014), the evidence of racial tensions we found in our analysis indicates otherwise. Despite a range of efforts to reduce racial discrimination, our study indicates the need for additional initiatives to promote racial harmony and resist the othering of racial minorities.
@@
Our results highlight ethical concerns about the potential ‘nudge effects’ of autocomplete suggestions on YouTube users. Graham (2023) noted that autocompletes can be biased in ways that might not be clear on an individual level but become apparent when evaluating a larger sample. Our analysis reveals evidence of similar aggregated discrimination in race-related queries on YouTube. For example, we find that while Black struggles to combat oppression are diminished, e.g., through the removal of references to “Black Lives Matter” in autocomplete suggestions, “White Lives Matter” continues to feature in these suggestions despite its known associations with hate groups. We also found traces of power imbalance through various ways of cultural appropriation reflected in YouTube autocompletes, showing how asymmetric racial values are embedded in YouTube’s technological configurations. Our unique methodological approach has helped us surface these biases, and it can also inform future analyses of autocomplete outputs, especially on social media and video-sharing platforms.
Prior studies have shown how biased autocomplete suggestions can impact users’ votes in political elections (Epstein et al., 2024), lead to hiring discrimination (Lin et al., 2024), and shift attitudes toward societal issues, e.g., support for fracking and legalizing the death penalty (Williams-Ceci et al., 2024). Thus, there is cause for concern that repeated exposure to racial stereotypes in YouTube autocompletes may adversely influence users’ attitudes and behaviors toward different racial groups. The black-box nature of such AI-based technologies is likely to prevent users from gaining access to information about how outputs were generated and how they could embed identity-based biases (Rai, 2020). In fact, users may be totally unaware of AI biases and fail to notice the influence of these biases on their attitudes (Williams-Ceci et al., 2024). Future work can build upon this research by analyzing the cognitive processes that mediate how YouTube users are influenced by biases in autogenerated content on the site.
Indeed, previous research on Google Search shows that its autocomplete outputs are sourced not just from web content and users’ search activity but also from external digitization projects, especially Google Books (Graham, 2023). Lack of information about how autocompletes are produced allows platforms to avoid taking the full ethical responsibility for problematic suggestions, even in cases when they result from machine learning biases rather than prior search queries. Therefore, we call for greater transparency regarding autocomplete sources and algorithms to facilitate a more precise attribution of problematic suggestions.
Our findings indicate the need for moderation strategies for the ethical deployment of AI-based tools such as autocompletes. Prior research on moderation in search engine autocompletes has shown that platforms moderate derogatory stereotypes about various demographic categories (Leidinger and Rogers, 2023). Google even has an autocompletion moderation policy to prevent autocompletes “that might be violent, sexually-explicit, hateful, disparaging or dangerous” (Sullivan, 2020). Yet, as our findings show, YouTube, owned by the same parent company as Google, perpetuates racial stereotypes, highlighting the need for improved policy and corporate-wide enforcement. As Rogers (2023) points out, platforms like Google should take actions that exceed mere ‘quiet patches’ of infamous autocompletions highlighted by journalists and researchers on English language sites by ensuring that adequate remedies occur for all their products and language offerings.
@@
This study has some limitations. We observed an under-representation of Hispanics in our analysis of autocompletes. However, it is possible that this result is influenced by other terms referring to the “Hispanic” race (e.g., “Latino,” “Mexican”) being used colloquially more often than similar counterparts (e.g., “Caucasian,” “African-American”) for other races. Future research on autocomplete audits should also collect queries using such colloquial terms.

### v3

@@ 4.5 Manner
Inappropriate Behavior. This category contains autocompletes that describe Black people as lazy or behaving inappropriately in social settings. One longstanding negative stereotype of Black people maintains that they are lazy, which is also posed as an explanation for other undesirable attributes, such as their disproportionate poverty (Gilens, 1996). The autocomplete “black people can’t hear smoke detectors” depicts this stereotype; it led to videos that show Black people as being too lazy to change the low battery of smoke detectors even though their alarms kept sounding. Figure 9 presents a series of YouTube shorts that show how creators promote this stereotype via memes.
@@ 5 Discussion
We began this study with the goal of examining the information that YouTube’s search feature provides to users seeking video content on race-related topics. As expected, we found the presence of racial biases and stereotypes in search autocomplete outputs, revealing the critical problems with search moderation. However, we also found that the participatory nature of YouTube has contributed to a more complex picture of race-based information in its search outputs, marked by desires for a race-neutral society, condemnations of historical oppressions, and curiosities about other races. We discuss these complexities below.
First, our analysis reveals that race-related autocompletes embody troubling stereotypes regarding Blacks’ personal hygiene, financial abilities, language use, aggression, and laziness. Many search outputs reinforce historical biases against Blacks (e.g., the poor Black, the ignorant Black, the criminal Black), show a lack of awareness about them, and perpetuate their otherization. While we found instances of negative stereotypes against Whites (e.g., regarding their athletic ability), they occurred less frequently than stereotypes against racial minorities. This asymmetry underscores the potential of search autocomplete feature to contribute to representational harm.
We found that both mainstream media and user-generated content contribute to the production of racial stereotypes in YouTube search outputs. Notably, we observed that YouTube search amplifies distorted racial views that appear in mainstream entertainment media by reproducing titles or dialogues from existing media content (e.g., “white men can’t jump”). On the other hand, amateur YouTube creators, seemingly driven by the platform’s commercial logic (Boxman-Shabtai, 2019), also contribute to racial stereotypes, often delivering them in the form of ethnic humor and memes. Curiously, we found that such ethnic humor is sometimes delivered by creators belonging to racial minorities (e.g., Steven He O’Byrne), presumably to gain an audience by propagating stereotypes about their own race with impunity.
Besides racial stereotypes, we also identified traces of interracial tensions through various ways of cultural appropriation, anxieties around interracial relationships, and concerns about persistent underrepresentation in the media. These insights reflect contemporary racial dynamics in the U.S. and suggest shifting terrains of racial discord, visibility, and exclusion that warrant further inquiry.
Balanced against the above problematic trends, we found that YouTube search outputs reproduce some creators’ attempts to dismantle racial stereotypes, often through deploying humor (e.g., “every black man should have a latte.”). However, such dismantling often occurred only after viewing the videos in the corresponding search results; in many cases, the autocomplete text alone—short, decontextualized, and hyper-visible—perpetuates racial biases (e.g., “black woman can’t buy a dress,” “white lives matter”). We observed discursive events that showed a repudiation of racial inequalities and hopes for a post-race society (e.g., “is there a hispanic disney princess”). We also found potential attempts at promoting positive representations of racial minorities (e.g., “black people can dance to anything”), seeking practical information about one’s own race, and genuine curiosities about other races. Additionally, we found depictions of struggles and inequalities faced by racial minorities (e.g., “asian men are bitter,” “can’t keep a black man down”). Thus, the user-generated content that drives YouTube serves to produce search outputs that go beyond racial stigmas and reveal the ideological messiness of the participatory culture (Khan and Malik, 2022).
@@
Our results highlight ethical concerns about the potential ‘nudge effects’ of autocomplete suggestions on YouTube users. Graham (2023) noted that autocompletes can be biased in ways that might not be clear on an individual level but become apparent when evaluating a larger sample. Our analysis reveals evidence of similar aggregated discrimination in race-related queries on YouTube. For example, we find that while Black struggles to combat oppression are diminished, e.g., through the removal of references to “Black Lives Matter” in autocomplete suggestions, “White Lives Matter” continues to feature in these suggestions despite its known associations with hate groups.
Examining the relative prevalence of different themes in our data, we note that skin tone, which serves as an explicit marker of race, appeared much less frequently than other categories like culture and ability in search autocompletes. This aligns with prior findings by Walters et al. (2024) on race-related Facebook memes, where explicit depictions of race and appeals to racism appeared much less frequently than implicit appeals to forward a racist agenda. This suggests a need to attend to traditionally palatable rhetorics of racism rather than just overtly racist speech on digital platforms.
Building on our findings, we recommend a careful development of moderation policies and practices for the ethical deployment of search autocomplete tools. As Rogers (2023) points out, search platforms should take actions that exceed mere ‘quiet patches’ of infamous autocompletions highlighted by journalists and researchers—they must institute overarching mechanisms that proactively detect and regulate problematic suggestions. Like many AI-based technologies (Rai, 2020), autocomplete feature does not inform users how it generates specific outputs and why they embed identity-based biases. Such lack of information allows search platforms to avoid taking the full ethical responsibility for problematic suggestions, even in cases when they result from machine learning biases rather than prior search queries or creators’ contributions. Therefore, we call for greater transparency regarding autocomplete sources and algorithms to facilitate a more precise attribution of problematic suggestions.
@@
This study has some limitations. We observed an under-representation of Hispanics in our analysis of autocompletes. However, it is possible that this result is influenced by other terms referring to the “Hispanic” race (e.g., “Latino,” “Mexican”) being used colloquially more often than similar counterparts (e.g., “Caucasian,” “African-American”) for other races. Future research on autocomplete audits should also collect queries using such colloquial terms. Our arguments about the influence of autocomplete biases on users would benefit from further engaging with the duality of how aggregated user behaviors and audience cues, in turn, feed into autocomplete outputs. While we attempted to obtain a temporally stable and geographically representative dataset, the YouTube platform constantly updates its search autocomplete suggestions, thus limiting the study’s generalizability. Although outside the scope of the current study, future research would benefit from a systematic comparison of search autocompletes (and their associated biases) on Google and YouTube sites and examining the reasons behind those differences.
6 Conclusion
This article has presented the first systematic analysis of racial stereotypes in YouTube’s search autocomplete outputs, addressing a major gap given YouTube’s cultural significance and its role as the second-largest search engine. Our thematic framework highlights asymmetric representation of different racial groups, evidence of harmful stereotypes, and hegemonic counter-struggles. We call for greater transparency and reform in content moderation to counteract identity-based harms enacted by search outputs.